Exercise 12-10 Skysong Industries has the following patents on its December 31, 2016, balance sheet. Patent Item Patent A Patent B Patent C Initial Cost $44,268 $17,040 $22,560 Date Acquired 3/1/13 7/1/14 9/1/15 Useful Life at Date Acquired 17 years 10 years 4 years The following events occurred during the year ended December 31, 2017. 1. Research and development costs of $247,000 were incurred during the year. 2. Patent D was purchased on July 1 for $37,848. This patent has a useful life of 91/2 years. 3. As a result of reduced demands for certain products protected by Patent B, a possible impairment of Patent Bs value may have occurred at December 31, 2017. The controller for Skysong estimates the expected future cash flows from Patent B will be as follows. Year 2018 2019 2020 Expected Future Cash Flows $2,200 2,200 2,200 The proper discount rate to be used for these flows is 8%. (Assume that the cash flows occur at the end of the year) Compute the total carrying amount of Skysong patents on its December 31, 2016, balance sheet. (Round answer to O decimal places, e.g. 8,564.) Total carrying amount$

media%2F233%2F233433e0-af4d-4c2a-88a0-5e

This question has two parts. Thank you.
Answer #1

Ans-1-Computing the total carrying amount of Skysong' patents on its December 31,2016 as follows:-

Patent Initial Cost Date Acquired

Useful Life

(In Years)

Useful Life

(In months)

Number of Months till December 31,2016 Amortization till December 31, 2016 Carrying Amount as on December 31, 2016
(A) (B) (C) (D)= (C) *12 (E) (F)=(A)*(E)/ (D) (G)=(A)-(F)
Patent A $44,268 3/1/2013 17 Years 204 46 $9,982 $34,286
Patent B $17,040 7/1/2014 10 Years 120 30 $4,260 $12,780
Patent C $22,560 9/1/2015 4 Years 48 16 $7,520 $15,040

2- Computing carrying amount of Skysong's Patents on its December 31, 2017:-

Patent Initial Cost Date Acquired

Useful Life

(In Years)

Useful Life

(In Months)

Number of Months Till December 31, 2017 Amortization till December 31, 2017 Carrying Amount as on December 31, 2017 before impairment Impairment Carrying Amount as on December 31, 2017
(A) (B) (C) (D)=(C)*12 (E) (F)=(A)*(E)/(D) (G)=(A)-(F) (H) (I)=(G)-(H)
Patent A $44,268 3/1/2013 17 Years 204 58 $12,586 $31,682 0 $31,682
Patent B $17,040 7/1/2014 10 Years 120 42 $5,964 $11,076 $5,407 $5,669
Patent C $22,560 9/1/2015 4 Years 48 28 $13,160 $9,400 0 $9,400
Patent D $37,848 7/1/2017 9.5 Years 114 6 $1,992 $35,856 0 $35,856

Working Note:-

Computing Discounted Future Cash Flows of patent B follows:-

Year Expected Cash Flow Discount Factor @8% Present Value
2018 $2,200 0.9259 $2,036.98
2019 $2200 0.8573 $1,886.06
2020 $2,200 0.7938 $1,746.36

$5,669.40

Impairment Loss of Patent B :-

Carrying Value of Patent B before Impairment= $11,076

Discounted Future Cash Flows of Patent B =$5,669

Impairment Loss= $11,076- $5,669

=$5,407